The area of a triangle is 37 square feet and the height is 10 feet. Find the base? I know that Area = 1/2 base height. The Area is 37 square feet and the height is 10 feet. I am lost can you help me? please?
Answer by arunpaul(104)

Area A = 37 sqft
height h = 10 ft
as u know area = (1/2)h*b=37 (where b = base)
or (1/2)10*b=37
b = (37*2/10)
b = 74/10=7.4 (Ans)
so base = 7.4ft